OpenStudy (anonymous):

Will give Medal! Which of the following is the equation of a circle whose center is at the origin and whose radius is 4? x(squared)+ y(squared) = 2 x(squared) + y(squared) = 4 x(squared) + y(squared) = 16

OpenStudy (anonymous):

THE GENERAL EQUATION OF A CIRCLE IS \[(x-a)^{2}+(y-b)^{2}=r ^{2}\] (a,b) is the coordinate of the centre and r is the radius since a and b is zero when centre lies at origini, it means that all the circles mentioned have centre at origin for radius to be 4, r^2=16 x^2+y^2=16 is the answer